A bicyclist increases her speed from 4.0 m/s to 6.0 m/s. The combined mass of the bicyclist and the bicycle is 55 kg. How much work did the bicyclist do in increasing her speed?




1
Expert's answer
2021-02-26T10:03:16-0500

Answer

The work is equivalent to change in kinetic energy

"E=W=\frac{m(v_f^2-v_i^2) }{2}"

"=\frac{55(6^2-4^2) }{2}\\=550Joule"